OpenStudy (anonymous):

Two particles each of mass 'm' are connected to a light weight string of length '2L' . a steady force 'F' is applied at the midpoint of the string (x=0) at a right angle to the initial position of the string. Show that the acceleration of each mass in the direction at 90 degrees to F is given by Ax = (Fx)/(2m((L^2) - (x^2) ))^(1/2)

OpenStudy (anonymous):

 force acting on one of the mass m:\[Fcos \theta =m*a\] where a is the accn of one of the mass along x axis. the reason behind writting cos theta is because, the force is acting along y-axis...but the particles move along y axis. the force acts along the string and the component of this force along the direction of displacement of the particle is fcos theta. \[Ftotal = 2Fcos \theta \] since there are two particles..... now force acting on 1 particle is :\[F= fcos \theta =(mLa/x)\] \[Ftotal= (2mLa/x)*\cos phi \] phi is the angle made by each component of force with the y axis. therefore, \[\cos \phi = (L ^{2}-x ^{2})^1/2 /L\] thus, \[Ftotal= 2fcos \theta = (2mL/x)*(L ^{2}-x ^{2})^1/2 /L\] this is equal to \[2m* a(total)\] therefore, \[a(total) = F/(2m) ^{2}*(L ^{2}-x ^{2})^{1/2} \]\[= F/\sqrt{(2m(L ^{2}}-x ^{2}))\]